Given the approximate pH of the following items, calculate the [H+]: a. Oranges 3.5 ?


1
Expert's answer
2021-05-09T00:01:08-0400

Solution:

pH = 3.5

We can convert between pH and [H+] using the following equation:

pH = - log[H+]

[H+] = 10-pH = 10-3.5 = 3.16×10-4

[H+] = 3.16×10-4 M


Answer: [H+] = 3.16×10-4 M